You draw two cards from a standard deck of cards (52 total cards), without replacing the first card before your second draw. What is the probability that both cards are either a king or a queen?
Found 2 solutions by Fombitz, dkppathak:
Answer by Fombitz(32388)
Answer by dkppathak(439)

You draw two cards from a standard deck of cards (52 total cards), without replacing the first card before your second draw. What is the probability that both cards are either a king or a queen?
there are four king and four queen in deck of 52 cards
probability for drawing 1 card from 52 card that is king =4/52
probability for drawing 1 card again with out replacing first card which is queen=4/51
probability for either king or queen will be =4/52+4/51
=4x51+4x52/52x51
=4(52+51)/52x51
=4x103/52x51
=412/2652
=0.1553
answer probability for either king or a queen =0.1553